40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Nov 11 22:11:15 2015, Koji, Update, SUS, PSL/IOO maintenance, TM SUS check up TM_SUS.pdfTM_SUS_SD.pdf
    Reply  Thu Nov 12 09:15:56 2015, Steve, Update, SUS, ITMX osems ITMX.png
       Reply  Fri Nov 13 15:20:24 2015, Steve, Update, SUS, OSEMs OSEMs.pngPOSbiasOFF.png
    Reply  Thu Nov 12 09:41:02 2015, Steve, Update, PSL, PMC PMC_ITMX.png
    Reply  Tue Nov 17 00:57:21 2015, ericq, Update, SUS, ITMX UL calmed? TM_nov17.pdfITMX_sensortrend.png
       Reply  Tue Nov 17 15:59:23 2015, rana, Update, SUS, ITMX UL calmed? 
Message ID: 11770     Entry time: Tue Nov 17 00:57:21 2015     In reply to: 11753     Reply to this: 11774
Author: ericq 
Type: Update 
Category: SUS 
Subject: ITMX UL calmed? 

After running dither alignment for all mirrors, all oplevs were recentered. (Except ETMY, since we did that earlier today.)

Looking at Koji's template for OSEM signals, the ITMX UL sensor noise floor seems more in line with the LL sensor, though there continues to be more noise than in other mirrors. 

 

Trending the sensor signals over the past 7 days, Koji's measurement looks to have been taken during a time when the UL sensor voltage had jumped down. Did someone squish the satellite box cable? I have not done so. 

I think that the step right at the end is due to a new POS offset of -2k counts, which I think Koji put into place earlier today. 

According to the wiki, the Vmax/2 values and the current values are: 

Sensor

Vmax/2

Vnow

UL

0.985

0.75

UR

0.855

0.865

LR

0.735

0.938

LL

0.775

0.60

SD

0.835

1.16

Attachment 1: TM_nov17.pdf  104 kB  | Hide | Hide all
TM_nov17.pdf
ELOG V3.1.3-